# PFGuard: A Generative Framework with Privacy and Fairness Safeguards

### v1

@@ 5.4 Analysis with Stronger Privacy, High-Dimensional Images
Table 4 shows the fairness and utility performances under these challenging conditions. We observe that DP generative models often exhibit extreme accuracy disparities even with a simplistic class bias setting, achieving over 90% accuracy for the majority class while achieving accuracy below 25% for the minority class. PFGuard consistently enhances the minority class performance and reduces accuracy disparity, while there is still room for improvements. Our results underscore the importance of tackling both privacy and fairness in future studies, encouraging more research in this critical area.
Table 4: Fairness and utility performances of private generative models with and without PFGuard on downstream tasks, evaluated on CelebA with binary class bias under $\varepsilon{=}1$. GS-WGAN is excluded due to lower image quality in this setting. Blue and red arrows indicate positive and negative changes, respectively. We provide the full table with standard deviations in Sec. E.4.
@@ 6 Related Work
We cover the private and fair data generation literature here and cover the 1) private-only data generation, 2) fair-only data generation, 3) privacy-fairness intersection literature in Sec. F. Compared to these lines of works, only a few works focus on private and fair data generation (Xu et al., 2021; Pujol et al., 2022). First, (Xu et al., 2021) proposes a two-step approach that removes bias from the training data via a fair pre-processing technique (Celis et al., 2020) and learns a DP generative model (Chanyaswad et al., 2019) from the debiased data. However, this framework is limited to low-dimensional structural data due to data binarization step in pre-precessing stage, which can incur significant information loss in high-dimensional image data. PFGuard, on the other hand, can generate high-dimensional image data with high quality. Second, (Pujol et al., 2022) proposes private data generation techniques satisfying causality-based fairness (Salimi et al., 2019), which consider the causal relationship between attributes. In comparison, PFGuard focuses on statistical fairness to achieve similar model performances for sensitive groups (Barocas et al., 2018). While causality-based approaches can better reveal the causes of discrimination than statistical approaches, modeling an underlying causal mechanism for real-world scenarios is also known to be challenging.
@@ 7 Conclusion
We believe our research addresses the critical issue of Trustworthy AI. Our focus on privacy and fairness underscores the need to design AI models that simultaneously safeguard individual privacy and mitigate biases without perpetuating them. In addition, our research and experiments are conducted with a strong commitment to ethical standards. All datasets used in this study, including publicly available human images, are widely used within the research community and do not contain sensitive or harmful content. Finally, we do note that choosing the right privacy and fairness measures for an application can be challenging and also depends on the social context. We also note that the use of multiple teacher does increase the cost of training, but provides more benefits particularly in balancing privacy and fairness.
@@
All datasets, methodologies, and experimental setups utilized in our study are described in detail in the supplementary material. More specifically, we provide a description of the proposed algorithm in Sec. C.1, details of datasets and preprocessing in Sec. D.1, and implementation details including hyperparameters in Sec. D.2 to ensure reproducibility.
@@ Appendix A Challenges of Extending Classification Techniques
Continuing from Sec. 3, we provide more details of potential challenges when one tries to extend fairness-privacy classification techniques (Jagielski et al., 2019; Mozannar et al., 2020; Tran et al., 2021b; 2022; Lowy et al., 2023) to generative settings due to the fundamentally different goals.
@@
Classification and generation settings often concentrate on different DP notions or rely on different assumptions, which hinders simple extensions of techniques between them. In the classification setting, Differential Privacy w.r.t. sensitive attribute (Jagielski et al., 2019) is commonly addressed (Jagielski et al., 2019; Mozannar et al., 2020; Tran et al., 2021b; 2022; Lowy et al., 2023), which considers the demographic group attribute as the only private information. This DP notion requires less DP noise compared to a more general notion of DP (Def. 2.1), which protects all input features, and enables a better privacy-utility tradeoff for DP classifiers. However, in the generative setting, a general notion of DP is mostly addressed, as presumed non-private features may in fact encode private information (e.g., pixel values in a facial image). Therefore, simply extending classification techniques to generative settings can be challenging, as it necessitates rigorous mathematical proofs for corresponding DP notions and may add a large DP noise when adapting to a general DP notion. Moreover, classification techniques can rely on convex objective functions (Tran et al., 2021a), but the assumption of convexity does not usually hold in generative models (Goodfellow et al., 2014).
@@
While recent studies have proposed fair variants of gradient sanitization (Xu et al., 2020; Esipova et al., 2022), directly adopting them in existing private generative models can undermine the original utility and privacy guarantee. To prevent aggressive gradient clipping in minority data groups, approaches to tune clipping threshold $C$ during training have been proposed, such as dynamically adjusting $C$ during training (Esipova et al., 2022) or utilizing different $C$ values w.r.t. groups (Xu et al., 2020). However, these adjustments of $C$ not only consume additional privacy budget, but also can significantly affect model utility, as private generative models often demonstrates high sensitivity in model convergence to these clipping values (Chen et al., 2020; Wang et al., 2021a; Dockhorn et al., 2022). Hence, given the limited privacy budget and the necessity to carefully set the value of $C$, these approaches of tuning $C$ may drastically change the original privacy-utility tradeoff of existing models to achieve fairness.
@@ Appendix B Differential Privacy
Continuing from Sec. 2 and Sec. 4.2, we provide more details on differential privcy (DP).
@@ B.1 Post-processing Property of Differential Privacy
Continuing from Sec. 2, we detail the post-processing property of DP. Let $G$ be an $(\varepsilon,\delta)$-DP generator, which produces the synthetic data from input random noise ${\mathbf{z}}\in\mathcal{Z}$. Then, the synthetic dataset $\tilde{D}=G({\mathbf{z}})$ is also satisfies $(\varepsilon,\delta)$-DP due to the post-processing theorem, as the random noise ${\mathbf{z}}$ is independent of the private dataset $D$, which is used to train the DP generator.
@@ B.2 Sensitivity Analysis of GNMax Aggregator
Given $\{T_{i}\}_{i=1}^{n_{T}}$ teachers, $c$ possible label classes, and an input data sample ${\mathbf{x}}$, the teachers’ vote count for the $j$-th class to an input ${\mathbf{x}}$ is denoted as:
$n_{j}({\mathbf{x}})=|\{i:T_{i}({\mathbf{x}})=j\}|\quad\text{for }j=1,...,c$ (7)
@@
$\mathbf{n}({\mathbf{x}})=(n_{1},\dots,n_{c})\in\mathbb{N}^{c}$ (8)
Since a single training data point only affects at most one teacher due to data disjointness, changing one data sample will at most change the votes by 1 for two classes, where we denote here as classes $i$ and $j$. Given the two adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ which differ by a single data point, let the aggregated vote counts are $\mathbf{n}=(n_{1},\dots,n_{c})$ and $\mathbf{n}^{\prime}=(n_{1}^{\prime},\dots,n_{c}^{\prime})$, respectively. The $l_{2}$-sensitivity (Def. 2.2) can be derived as follows:
$\Delta^{2}$ $=\max_{D,D^{\prime}}\|(n_{1},\dots,n_{c})-(n_{1}^{\prime},\dots,n_{c}^{\prime})\|_{2}$ (9)
@@ Appendix C PFGuard Framework
Continuing from Sec. 4.2, we provide more details on the PFGuard framework.
C.1 Training Algorithm
@@
Continuing from Sec. 4.2, we provide the pseudocode to describe the full training algorithm when PFGuard is integrated on top of a PTEL-based generative model.
@@
Input Training dataset $D$, ensemble of teacher model $T=\{T_{i}\}_{i=1}^{n_{T}}$ with each parameters $\theta_{T_{i}}$, batch size $B$, PTEL algorithm $\mathcal{M}({\mathbf{x}};T)$, teacher loss function $\mathcal{L}_{T}$, generator loss function
@@
1:Divide the dataset $D$ into subsets $\{D_{i}\}_{i=1}^{n_{T}}$
@@
5:         Draw a minibatch $\{{\mathbf{x}}_{i}\}_{i=1}^{B}\subseteq D_{i}$ with sampling ratio $w({\mathbf{x}})\propto h({\mathbf{x}})$ using Eq. 5
@@
7:         Update teacher model $T_{i}$ with $\mathcal{L_{T}}(\theta_{T};{\mathbf{x}},G({\mathbf{z}};\theta_{G}))$
@@
10:     Aggregate teacher output $\tilde{o}\leftarrow\mathcal{M}({\mathbf{x}};T)$
11:     Update generator model $G$ with $\mathcal{L}_{G}(\theta_{G};{\mathbf{z}},\tilde{o})$
12:     Accumulate privacy cost $\varepsilon$
13:end for
14:return Generator G, privacy cost $\varepsilon$
Notably, PFGuard requires to only modify the minibatch sampling process for training the teacher models (e.g., line 5 in the above pseudocode) to enable fair training within the training algorithm of PTEL-based generative models. PFGuard also preserves their privacy analyses (e.g., privacy cost $\varepsilon$ of each training epoch), as long as these models rely on data disjointness to derive sensitivity.
C.2 Setting Number of Teachers without Sensitive Attribute Labels
Continuing from Sec. 4.2, we discuss how to extend the proposed upper bound on the number of teachers (i.e., $\lfloor|D|\min_{s\in\mathcal{S}}p_{\text{bias}}(s)\rfloor$) in settings where the label of sensitive attribute ${\mathbf{s}}$ is unavailable. The proposed upper bound does not rely on full knowledge of $p_{\text{bias}}$, but the distribution w.r.t. sensitive attributes. Thus, given the training data, we can estimate the subgroup distributions using traditional techniques like K-means clustering (Macqueen, 1967) or random subset labeling (Forestier & Wemmert, 2016). We note that these estimations can be effective, but may introduce some errors or additional overhead, such as increased computational time.

### v2

@@ 5.4 Analysis with Stronger Privacy, High-Dimensional Images
Table 4 shows the fairness and utility performances under these challenging conditions. We observe that DP generative models often exhibit extreme accuracy disparities even with a simplistic class bias setting, achieving over 90% accuracy for the majority class while achieving accuracy below 25% for the minority class. PFGuard consistently enhances the minority class performance and reduces accuracy disparity, while there is still room for improvement. Our results underscore the importance of tackling both privacy and fairness in future studies, encouraging more research in this critical area.
Table 4: Fairness and utility performances of private generative models with and without PFGuard on downstream tasks, evaluated on CelebA with binary class bias under $\varepsilon{=}1$. GS-WGAN is excluded due to lower image quality in this setting. Blue and red arrows indicate positive and negative changes, respectively. We provide the full results with standard deviations in Sec. E.8.
@@ 6 Related Work
We cover the private and fair data generation literature here and cover the 1) private-only data generation, 2) fair-only data generation, 3) privacy-fairness intersection literature in Sec. F. Compared to these lines of works, only a few works focus on private and fair data generation (Xu et al., 2021; Pujol et al., 2022). First, (Xu et al., 2021) proposes a two-step approach that 1) removes bias from the training data via a fair pre-processing technique (Celis et al., 2020) and 2) learns a DP generative model (Chanyaswad et al., 2019) from the debiased data. However, this framework is limited to low-dimensional structural data due to data binarization step in pre-precessing stage, which can incur significant information loss in high-dimensional image data. PFGuard, on the other hand, can generate high-dimensional image data with high quality. Second, (Pujol et al., 2022) proposes private data generation techniques satisfying causality-based fairness (Salimi et al., 2019), which consider the causal relationship between attributes. In comparison, PFGuard focuses on statistical fairness to achieve similar model performances for sensitive groups (Barocas et al., 2018). While causality-based approaches can better reveal the causes of discrimination than statistical approaches, modeling an underlying causal mechanism for real-world scenarios is also known to be challenging.
@@ 7 Conclusion
We believe our research addresses the critical issue of Trustworthy AI. Our focus on privacy and fairness underscores the need to design AI models that simultaneously safeguard individual privacy and mitigate biases in training data. In addition, our research and experiments are conducted with a strong commitment to ethical standards. All datasets used in this study, including publicly available human images, are widely used within the research community and do not contain sensitive or harmful content. There are also limitations, and we note that choosing the right privacy and fairness measures for an application can be challenging and also depends on the social context. We also note that the use of multiple teacher does increase the cost of training, but can provide more benefits particularly in balancing privacy and fairness.
@@
All datasets, methodologies, and experimental setups used in our study are described in detail in the supplementary material. More specifically, we provide a description of the proposed algorithm in Sec. C.2, details of datasets and preprocessing in Sec. D.1, and implementation details including hyperparameters in Sec. D.2 to ensure reproducibility.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00509258, Global AI Frontier Lab). This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(No. RS-2022-NR070121).
@@ Appendix A Challenges of Extending Classification Techniques
Continuing from Sec. 3, we provide more details of potential challenges when one tries to extend fairness-privacy classification techniques (Jagielski et al., 2019; Mozannar et al., 2020; Tran et al., 2021b; 2022) to generative settings due to the fundamentally different goals.
@@
Classification and generation settings often concentrate on different DP notions or rely on different assumptions, which hinders simple extensions of techniques between them. In the classification setting, Differential Privacy w.r.t. sensitive attribute (Jagielski et al., 2019) is commonly addressed (Jagielski et al., 2019; Mozannar et al., 2020; Tran et al., 2021b; 2022), which considers the demographic group attribute as the only private information. This DP notion requires less DP noise compared to a more general notion of DP (Def. 2.1), which protects all input features, and enables a better privacy-utility tradeoff for DP classifiers. However, in the generative setting, a general notion of DP is mostly addressed, as presumed non-private features may in fact encode private information (e.g., pixel values in a facial image). Therefore, simply extending classification techniques to generative settings can be challenging, as it necessitates rigorous mathematical proofs for corresponding DP notions and may add a large DP noise when adapting to a general DP notion. Moreover, classification techniques can rely on convex objective functions (Tran et al., 2021a), but the assumption of convexity does not usually hold in generative models (Goodfellow et al., 2014).
@@
While recent studies have proposed fair variants of DP-SGD (Xu et al., 2020; Esipova et al., 2022), directly adopting them in existing private generative models can undermine the original utility and privacy guarantee. To prevent aggressive gradient clipping in minority data groups, approaches to tune clipping threshold $C$ during training have been proposed, such as dynamically adjusting $C$ during training (Esipova et al., 2022) or utilizing different $C$ values w.r.t. groups (Xu et al., 2020). However, these adjustments of $C$ not only consume additional privacy budget, but also can significantly affect model utility, as private generative models often demonstrate high sensitivity in model convergence to these clipping values (Chen et al., 2020; Wang et al., 2021a; Dockhorn et al., 2022). Hence, given the limited privacy budget and the necessity to carefully set the value of $C$, these approaches of tuning $C$ may drastically change the original privacy-utility tradeoff of existing models to achieve fairness.
Here, we do not claim that extending classification techniques to generative settings is always impossible, but introduce the challenges that can complicate such extensions. To exemplify some possible cases, we extend classification methods from Esipova et al. (2022) and Xu et al. (2020) to generative settings, using them as baselines in our experiments – see Sec. 5 for results.
@@ Appendix B Differential Privacy
Continuing from Sec. 2 and Sec. 4.2, we provide more details on Differential Privacy (DP).
@@ B.1 Post-processing Property of Differential Privacy
Continuing from Sec. 2, we detail the post-processing property of DP as defined below.
@@
Due to the above post-processing theorem, a synthetic dataset $\tilde{\mathcal{D}}=G({\mathbf{z}})$ generated from an $(\varepsilon,\delta)$-DP generator $G$ with random noise ${\mathbf{z}}\in\mathcal{Z}$ also satisfies $(\varepsilon,\delta)$-DP, as the random noise ${\mathbf{z}}$ is independent of the private training dataset used for the DP generator.
@@ B.2 Sensitivity Analysis of GNMax Aggregator
Given $\{T_{i}\}_{i=1}^{n_{T}}$ teachers, $c$ possible label classes, and a query input $\bar{{\mathbf{x}}}$, the teachers’ vote counts for the $j$-th class to a query input $\bar{{\mathbf{x}}}$ is denoted as:
$n_{j}(\bar{{\mathbf{x}}})=|\{i:T_{i}(\bar{{\mathbf{x}}})=j\}|\quad\text{for }j=1,...,c$ (7)
@@
$\mathbf{n}(\bar{{\mathbf{x}}})=(n_{1},\dots,n_{c})\in\mathbb{N}^{c}$ (8)
Since a single training data point only affects at most one teacher due to data disjointness, changing one data sample will at most change the votes by 1 for two classes, where we denote here as classes $i$ and $j$ without loss of generality. Given the two adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ which differ by a single data point, let the aggregated vote counts be $\mathbf{n}=(n_{1},\dots,n_{c})$ and $\mathbf{n}^{\prime}=(n_{1}^{\prime},\dots,n_{c}^{\prime})$, respectively. The $l_{2}$-sensitivity (Def. 2.2) can be derived as follows:
$\Delta^{2}$ $=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|(n_{1},\dots,n_{c})-(n_{1}^{\prime},\dots,n_{c}^{\prime})\|_{2}$ (9)
@@ Appendix C PFGuard Framework
Continuing from Sec. 4.2, we provide more details on the proposed PFGuard framework.
C.1 Privacy anlaysis of PFGuard
Continuing from Sec. 4.2, we provide a theoretical proof of how PFGuard preserve the sensitivity of an arbitrary PTEL mechanism, and thus the sensitivity of an arbitrary PTEL-based generative model.
PTEL-based Generative Models
We first characterize the common training scheme of PTEL-based generative models (Jordon et al., 2018; Chen et al., 2020; Long et al., 2021; Wang et al., 2021a). To train a DP generator $G$ parametrized with $\theta_{G}$, the goal is to make its training algorithm $\mathcal{A}:\mathcal{D}\rightarrow G$ – which inputs the training dataset $\mathcal{D}$ and outputs the generator $G$ – satisfy DP (Long et al., 2021). Given the typical training scheme of stochastic gradient descent, which updates $\theta_{G}$ with gradient information $\bm{g}_{G}^{\text{up}}$, the training algorithm satisfies DP if $\bm{g}_{G}^{\text{up}}$ is privatized with a DP mechanism due to the post-processing property of DP (Thm. B.1(Abadi et al., 2016). To produce privatized gradient $\tilde{\bm{g}}_{G}^{\text{up}}$, PTEL-based generative models typically proceed in two steps as follows:
  • (1)
    Training teacher ensemble from disjoint data subsets. The private dataset $\mathcal{D}$ is first divided into $n_{T}$ disjoint subsets $\{\mathcal{D}_{i}\}_{i=1}^{n_{T}}$, where each subset $\mathcal{D}_{i}$ is uniquely used to train a teacher model $T_{i}$ with its parameter $\theta_{T_{i}}$. Thus, a total of $n_{T}$ teachers are trained and form a teacher ensemble $T=\{T_{i}\}_{i=1}^{n_{T}}$.
  • (2)
    Training generator by querying teacher ensemble. The target generator $G$ with the parameter $\theta_{G}$ is trained by interacting solely with the teacher ensemble $T$, without accessing the original dataset $\mathcal{D}$. Given a random input noise ${\mathbf{z}}$, the generator $G$ generates an output $G({\mathbf{z}})$ and queries the teacher ensemble $T$ with $G({\mathbf{z}})$ for their supervision. Each teacher in $T$ votes on the gradient $\bm{g}_{G}^{\text{up}}$ for the query input $G({\mathbf{z}})$, resulting in a vote count $\mathbf{n}(G({\mathbf{z}}))$. This vote count is processed by the PTEL mechanism $\mathcal{M}$ (e.g., GNMax aggregator in Sec. 4.2) to finally produce a DP-sanitized gradient $\tilde{\bm{g}}_{G}^{\text{up}}$. The generator $G$ updates its parameter $\theta_{G}$ using $\tilde{\bm{g}}_{G}^{\text{up}}$.
Thus, given a PTEL mechanism $\mathcal{M}$ with $(\varepsilon,\delta)$-DP guarantee for each iteration and a total of $N$ training iterations, the total DP guarantee of the training algorithm $\mathcal{A}$ can be computed via the composition theorem of DP (Dwork et al., 2006b).
Sensitivity Preservation
We now provide a theoretical proof of how training with PFGuard preserves the original privacy analysis of an arbitrary PTEL-based generative model $G$ as long as data disjointness is enforced. As explained above, the privacy analysis of the PTEL-based generative model depends on the 1) $(\varepsilon,\delta)$-DP guarantee of the given PTEL mechanism $\mathcal{M}$ for each training iteration and the 2) total number of training iterations $N$. Here, we focus more on the theoretical DP guarantee than the practical DP guarantee, so we assume that $N$ is preserved with the fair sampling algorithm ${Sample}(\cdot)$ used in PFGuard. That means, if $\mathcal{M}$ and $\mathcal{M}\circ\verb||{Sample}$ – PTEL mechanisms without and with PFGuard – result in the same $(\varepsilon,\delta)$-DP guarantee for each iteration, the same total DP guarantee is preserved. In particular, we show how $\mathcal{M}$ and $\mathcal{M}\circ\verb||{Sample}$ result in the same sensitivity (Def. 2.2), which ensures the same $(\varepsilon,\delta)$ values given the same amount of DP noise.
Since the PTEL mechanism $\mathcal{M}$ operates on a vote count $\mathbf{n}(\bar{{\mathbf{x}}})$ from the teacher ensemble $T$ given a generated sample as a query input (i.e., $\bar{{\mathbf{x}}}=G({\mathbf{z}})$), we can denote $\mathbf{n}(\bar{{\mathbf{x}}})$ = $\mathbf{n}(\bar{{\mathbf{x}}};T)$. Let $\mathcal{A}_{T}$ be the training algorithm of $T$, where $\mathcal{A}_{T}(\mathcal{D}_{i})=T_{i}$ and $\mathcal{A}_{T}(\mathcal{D})=T=\{T_{i}\}_{i=1}^{n_{T}}$ due to data disjointness used in training of $n_{T}$ teachers. Since ${Sample}(\cdot)$ – the proposed minibatch sampling algorithm – operates on $\mathcal{A}_{T}$ and samples a data subset $\mathcal{B}_{i}\subseteq\mathcal{D}_{i}$, data disjointness is preserved in $\mathcal{B}_{i}$. We thus can derive $\mathcal{A}_{T}(\verb||{Sample}(\mathcal{D}_{i}))=\mathcal{A}_{T}(\mathcal{B}_{i})=T_{s,i}$ and $\mathcal{A}_{T}(\verb||{Sample}(\mathcal{D}))=T_{s}=\{T_{s,i}\}_{i=1}^{n_{T}}$, resulting in a different teacher ensemble due to changed minibatches.
Let the original sensitivity value over $\mathbf{n}(\bar{{\mathbf{x}}};T)=\mathbf{n}(\bar{{\mathbf{x}}};\mathcal{A}_{T}(\mathcal{D}))$ be $k$. Then, due to the definition of the sensitivity, the following holds:
$\Delta^{2}$ $=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|\mathbf{n}(\bar{{\mathbf{x}}};\mathcal{A}_{T}(\mathcal{D}))-\mathbf{n}(\bar{{\mathbf{x}}};\mathcal{A}_{T}(\mathcal{D}^{\prime}))\|_{2}$ (12)
$=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|\mathbf{n}(\bar{{\mathbf{x}}};T_{1},...,T_{n_{T}-1},T_{n_{T}})-\mathbf{n}(\bar{{\mathbf{x}}};T_{1},...,T_{n_{T}-1},T^{\prime}_{n_{T}}))\|_{2}$ (13)
$=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|(n_{1},\dots,n_{c})-(n_{1}^{\prime},\dots,n_{c}^{\prime})\|_{2}$ (14)
$=\max_{n_{i},n_{i}^{\prime},n_{j},n_{j}^{\prime}}\sqrt{(n_{i}-n_{i}^{\prime})^{2}+(n_{j}-n_{j}^{\prime})^{2}}\leq k$ (16)
where the second equality follows from 1) adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ differing by a single data sample and 2) data disjointness, where a single data sample can affect a particular teacher that receives the data partition including the data sample); we denote $T_{n_{T}}$ as the affected teacher without loss of generality. The last equality denotes how much one teacher can maximally contribute to the PTEL voting scheme, captured by the sensitivity value.
We now compute the sensitivity value over $\mathbf{n}(\bar{{\mathbf{x}}};T_{s})=\mathbf{n}(\bar{{\mathbf{x}}};\mathcal{A}_{T}(\verb||{Sample}(\mathcal{D})))$ as follows:
$\Delta^{2}$ $=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|\mathbf{n}(\bar{{\mathbf{x}}};\mathcal{A}_{T}(Sample(\mathcal{D})))-\mathbf{n}(\bar{{\mathbf{x}}};\mathcal{A}_{T}(Sample(\mathcal{D}^{\prime})))\|_{2}$ (17)
$=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|\mathbf{n}(\bar{{\mathbf{x}}};T_{s,1},...,T_{s,n_{T}-1},T_{s,n_{T}})-\mathbf{n}(\bar{{\mathbf{x}}};T_{s,1},...,T_{s,n_{T}-1},T^{\prime}_{s,n_{T}}))\|_{2}$ (18)
$=\max_{\mathcal{D},\mathcal{D}^{\prime}}\|(v_{1},\dots,v_{c})-(v_{1}^{\prime},\dots,v_{c}^{\prime})\|_{2}$ (19)
$=\max_{v_{i},v_{i}^{\prime},v_{j},v_{j}^{\prime}}\|(0,\dots,0,v_{i}-v_{i}^{\prime},0,\dots,0,v_{j}-v_{j}^{\prime},0,\dots,0)\|_{2}$ (20)
$=\max_{v_{i},v_{i}^{\prime},v_{j},v_{j}^{\prime}}\sqrt{(v_{i}-v_{i}^{\prime})^{2}+(v_{j}-v_{j}^{\prime})^{2}}\leq k$ (21)
Here, the second equality follows similarly from adjacent datasets and data disjointness, where we denote $T_{s,n_{T}}$ as the affected teacher. Also, the last equality again captures how much one teacher can maximally contribute, which is a property of PTEL’s voting scheme and is independent of the sampling algorithm – thus being k. As a result, using the sampling technique in PFGuard preserves the sensitivity value as long as the PTEL mechanism enforces data disjointness, although the sampling results in a different teacher ensemble compared to the original PTEL-based generative model. We note that this sensitivity analysis remains valid when $\mathcal{B}$ includes duplicate samples due to potential oversampling because it still affects one teacher due to data disjointness.
C.2 Training Algorithm
@@
Continuing from Sec. 4.2, we provide the pseudocode describing the full training algorithm of PFGuard when integrated into a PTEL-based generative model.
@@
Input Training dataset $\mathcal{D}$, ensemble of teacher model $T=\{T_{i}\}_{i=1}^{n_{T}}$ with each parameters $\theta_{T_{i}}$, batch size $B$, PTEL mechanism $\mathcal{M}({\mathbf{x}};T)$, teacher loss function $\mathcal{L}_{T}$, generator loss function
@@
1:Divide the dataset $\mathcal{D}$ into subsets $\{\mathcal{D}_{i}\}_{i=1}^{n_{T}}$
@@
5:         Draw a minibatch $\{{\mathbf{x}}_{i}\}_{i=1}^{B}\subseteq\mathcal{D}_{i}$ with sampling ratio $w({\mathbf{x}})\propto h({\mathbf{x}})$ using Eq. 5
@@
7:         Update teacher model $T_{i}$ with $\mathcal{L_{T}}(\theta_{T_{i}};{\mathbf{x}},G({\mathbf{z}};\theta_{G}))$
@@
10:     Draw a set of random noise $\{{\mathbf{z}}_{i}\}_{i=1}^{B}$ from input random noise distribution $p_{z}$ of $G$
11:     Generate synthetic data samples $G({\mathbf{z}};\theta_{G})$
12:     Aggregate teacher output with PTEL mechanism $\tilde{o}\leftarrow\mathcal{M}(G({\mathbf{z}};\theta_{G});T)$
13:       where the voting is on gradient of $\mathcal{L}_{G}(\theta_{G})$
14:     Update generator model $G$ with $\tilde{o}$
15:     Accumulate privacy cost $\varepsilon$
16:end for
17:return Generator $G$, privacy cost $\varepsilon$
As shown in Algorithm C.2, PFGuard requires only a modification to the minibatch sampling process during teacher model training to achieve fairness (Line 5). Despite this modification, PFGuard preserves the original privacy guarantee of the given PTEL-based generative model, as long as its sensitivity analysis relies on data disjointness – see Sec. C.1 for more details.
C.3 Extension
Continuing from Sec. 4.2, we provide more details on the extensibility of PFGuard.
Setting Number of Teachers without Sensitive Attribute Labels
We discuss how to extend the proposed upper bound on the number of teachers (i.e., $\lfloor|\mathcal{D}|\min_{s\in\mathcal{S}}p_{\text{bias}}(s)\rfloor$) when the sensitive attribute label ${\mathbf{s}}$ is unavailable. Note that the proposed bound does not rely on full knowledge of $p_{\text{bias}}({\mathbf{x}})$ for data points to be generated (e.g., images), but instead relies on the distribution $p_{\text{bias}}({\mathbf{s}})$ w.r.t. sensitive attributes. Thus, we can estimate the sensitive attribute distribution in the training data using traditional techniques such as K-means clustering (Macqueen, 1967) or random subset labeling (Forestier & Wemmert, 2016) and use the estimated distribution in the proposed bound. We note that while these estimations can be effective, they may introduce errors or additional overhead, such as increased computational time.
Integration with Additional Fairness Techniques
PFGuard supports additional integrations of existing fairness methods if they meet two conditions: 1) they apply exclusively to teacher models, ensuring no direct impact on the target generator, and 2) they maintain data disjointness – each sample affects only one teacher – which is a fundamental requirement for PTEL’s privacy guarantees (see Sec. 4.2 for details). For example, GOLD (Mo et al., 2019), a fairness method designed to support Rawlsian Max-Min fairness (Joseph et al., 2016), is compatible with PFGuard. GOLD computes log density ratio estimates to identify worst-group samples and reweight the discriminator (i.e., teacher) loss to improve performance on these samples to achieve Rawlsian Max-Min fairness. Since this reweighting is applied solely to teacher models and preserves data disjointness, PFGuard with GOLD maintains the original privacy guarantees of the underlying PTEL generator – see privacy-fairness-utility tradeoff when additionally applying GOLD in Sec. E.6.
Synergy with Adversarial Training
While PFGuard’s training scheme is loss-agnostic and does not necessarily require adversarial training, PFGuard can be particularly synergistic with adversarial training. We discuss with two reasons: (1) the min-max optimization and (2) the assumption of an optimal discriminator in adversarial training. First, adversarial learning optimizes conflicting goals (i.e., the min-max optimization) using the two components of the generator and the discriminator. Since privacy and fairness can also have conflicting goals, assigning them to separate components may better balance these goals. Additionally, since the convergence of adversarial training depends heavily on the discriminator’s optimality (Goodfellow et al., 2014), assigning fairness specifically to the discriminator may effectively guide the generator toward an optimal state that is also fair.